• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jroche

    Log entry time 16:17:50 on July 15, 2005

    Entry number 148077

    keyword=day shift summary

    No target = no beam

    We started the shift by going to restricted access. The goal was to cooldown the target and to finish recovering the HRS after the power outage from yesterday.

    The HRS are now ready to go and on stand by at 0.5 GeV. The scanners are fully operational.

    The target cooldown did not happen. From what I understand, things are still being sorted out as I write. I'm sure more will be logged on this issue later on.